OpenStudy (anonymous):

if the factors of the polynomial are x-3 and x-4 what value of x makes the polynomial 0?

OpenStudy (raden):

solve for x's x-3 = 0 or x-4 = 0

OpenStudy (anonymous):

wait what

OpenStudy (raden):

x-3=0 can you solve for x

OpenStudy (raden):

also x-4=0 solve for x

OpenStudy (anonymous):

um no?

OpenStudy (raden):

x-3=0 both sides add by 3, giving us x-3+3 = 0 + 3 x = 3 it is the 1st do same idea to solve x-4 = 0